A standard deck of cards contains 52 cards. One card is selected from the deck.

​(a)

Compute the probability of randomly selecting a two or seven.

​(b)

Compute the probability of randomly selecting a two or seven or ace.

​(c)

Compute the probability of randomly selecting a six or diamond.

​(Type an integer or a decimal rounded to three decimal places as​ needed.)
Answer #1

a)

P(Two or Seven) = P(Two ) + P(seven)

= 4/52 + 4/52

= 8/52

= 0.154

b)

P(two or seven or ace) = P(two) + P(seven) + P(ace)

= 4/52 + 4/52 + 4/52

= 12/52

= 0.231

c)

P(six or diamond) = P(sex) + P(diamond) - P(sex of diamond)

= 4/52 + 13/52 - 1/52

= 16/52

= 0.308